# MIRI spectrophotometry of GN-z11: Detection and nature of an optical red continuum component

### v1

@@ 4.3 SF + AGN SED modelling
In addition to the young starburst (Bunker+23; Tacchella+23_GNz11; Alvarez-marquez+25), some studies indicates the presence of high critical density lines and broad emission lines in the ultraviolet suggesting that GN-z11 could host a type 1 active black hole (Maiolino+24_BH). To evaluate the presence of a potential AGN in GN-z11, we also consider the combination of star-formation and AGN (Type 1 or 2). The potential presence of an AGN would naturally produce a red component associated to the hot dust emission from the torus around an accreting black hole. The best-fit SED models for the different SF+AGN scenarios are presented in the top panel of Fig. 5. Specifically, CIGALE results including a type 1 and 2 AGNs are represented in blue and green, respectively, while SYNTHESIZER-AGN result is displayed as a red line. Tables 2 and 3 summarise the photometric and physical results of these models.
@@
Adding a type 1 AGN to the mixed stellar population modelled by CIGALE does not significantly improve the SED fitting beyond 0.66 m. Top and middle panels from Fig. 5 show how the type 1 AGN starts to dominate the emission only above 0.9 m and the best-fit modelled F770W and F1000W fluxes are well below the observed values by factors 1.5 and 2, respectively. The type 1 AGN is modelled in SYNTHESIZER-AGN by a QSO composite template (see Sect. 3.3) which includes the contribution from broad emission lines and the dust emission in the optical-red and near-infrared spectral range. Top and bottom panels in Fig. 5 shows that the QSO template is almost flat until rest-frame 1 m and it is not able to reproduce the observed fluxes in the MIRI reddest filters. Only when including an ad-hoc extra dusty torus (i.e., highly obscured QSO, see Sect. 3.3) the modelled F770W and F1000W fluxes (257 and 258 nJy, respectively) are compatible with the observed ones. However, the addition of this extra hot dust emission to the empirical spectrum of the QSO is not physically justified. As shown in Figure 3, the 0.7 m emission from the empirical 1 QSO SED agrees with the spectrum of the = 7.1 luminous QSO J1120+0641, for which JWST/MIRI spectroscopy has revealed a dust emission in the dusty torus with a temperature of about 1400 K (Bosman2024). This agreement in the SED between the low-intermediate- and high- QSOs difficults the physical justification for the presence of such an extra dusty torus emission.
@@
We consider next the presence of a type 2 AGN, whose UV and optical emission is expected to be obscured by the torus. According to the CIGALE best-fit model, the steep optical and near-IR continuum from the type 2 AGN emission increases the SED modelled fluxes up to 254 nJy for the F770W filter, being compatible with the MIRI measured fluxes (see Table 2). Despite the F1000W photometry not being taken into account during the SED fitting due to its low significance, the type 2 scenario yields a F1000W flux (285 nJy) compatible within the uncertainties, whereas the SED best-fit lies below the upper limit for F2100W. Although the available AGN models in CIGALE do not include the contribution from the emission lines produced by the AGN, the analysis carried out indicates that assuming a type 2 AGN yields a better agreement with the observed MIRI fluxes (i.e., optical to red continuum SED slope) than using different type 1 scenarios.
@@
Regardless of the AGN type considered, we derive similar ages for the different stellar populations. Specifically, we estimate ages of 4 6 and 20 40 Myr for the young and mature stellar populations, respectively. Ages for the stellar populations have been derived independently for the spatially resolved point-source and extended stellar components ( 11 and 35 Myr, respectively) identified in NIRCam images (Tacchella+23_GNz11). Although the resolution of the MIRI images (i.e., FWHM = 0.27 for F770W) does not allow us to spatially resolve these two components (rest-UV = 0.05 for the extended one), our derived ages are compatible the previous NIRCam photometry.
@@
To ensure the robustness of these results, we replicated the CIGALE SED fits using the SKIRTOR AGN models (Stalevski+12; Stalevski+16), which include a clumpy dust distribution for the torus. These AGN models yield similar results, where only a type 2 AGN can explain the F770W and F1000W fluxes. In next section, we will discuss the general physical scenario drawn from this SED best-fit along with other non-AGN possibilities.
@@ 4.4 Origin of the red excess emission
As already shown in the previous sections, the flux in the F770W filter, and in the low-significance (3) F1000W value, shows an excess that cannot be fully explained by the combination of a young and intermediate age stellar populations, even when including the contribution by the nebular emission (i.e., the narrow H line detected with the MRS). This excess hints to the presence of a source radiating at (rest-) wavelengths of 0.66 to 0.86 m. The excess at these wavelengths can be interpreted as the emission produced by either the presence of a dusty torus of an AGN or dust associated with the extremely compact starburst identified in GN-z11.
@@ 4.4.1 Dust in AGN: Dusty torus emission
Although GN-z11 has not been detected in X-rays, a 3 upper limit of (210 keV) ¡ 3 1043 erg s-1 has been derived assuming the typical photon index for narrow-line Seyfert 1 galaxies (Maiolino+24_BH). In addition, a value 2.9 1043 erg s-1, based on its H luminosity (Ho+01; Alvarez-marquez+25), could also be consistent with hosting a type 2 AGN. These upper limits agree well with the predicted (210 keV) based on the SED-fitting assuming the presence of type 1 and 2 AGNs. Using the X-ray extension of the CIGALE models (X-CIGALE, Yang+20), we obtain a (210 keV) = (0.9 0.1) and (2.1 0.3) 1043 erg s-1 for the best-fit SED considering the type 1 and 2 AGNs, respectively. Similar X-ray 2-10 keV luminosities (i.e., 1.8 1043 and 2.6 1043 erg s-1, respectively) are obtained when applying the (210 keV) relation from Marconi+04 and the bolometric luminosity derived from the SED fitting (see Table 3). In summary, the predicted X-ray luminosities of the best-fit stellar plus AGN (both type 1 and 2) SEDs are compatible with the no detection of X-ray emission in GN-z11.
@@
However, the scenario of a type 1 AGN is not favoured by the present results. The potential presence of an undetected, weak broad H line of 2 3 10-18 erg s-1 cm-2 tracing a type 1 AGN (Alvarez-marquez+25) would contribute to about 18 27 nJy to the F770W flux, not sufficient to explain the excess of 91 nJy found during the SED fitting (see Sect. 4.3). In addition, Figure 3 displays how luminous intermediate 1 (Glikman+06; Selsing+16) and 7 QSOs (Bosman2024) have a spectral energy distribution at rest-wavelengths 0.6 0.8 m flatter than otherwise detected at F770W and F1000W in GN-z11. The SED of these QSOs are consistent with hot dust from the black hole surrounding torus emitting at a temperature of about 1400 K, contributing at wavelengths of 1 m and above (e.g., Bosman2024). Moreover, our SYNTHESIZER-AGN best SED fit combining the stellar populations and type 1 AGN requires an additional ad-hoc red source (see Sect 4.3) in addition to the hot ( 1300 1400 K) dust component to explain the observed flux excess at 0.66 and 0.86 m. As already mentioned in the previous section, this extra component is not physically motivated, and therefore the scenario of a standard type 1 AGN is not supported by the present data.
@@
Under the scenario of a Seyfert 2 contribution, the AGN bolometric luminosity derived by CIGALE can be used to estimate the size and mass of the dusty torus assuming its physical parameters (i.e., radial extend and inclination of the torus, number of clouds along the equator of the torus, angular and radial distribution of clouds, optical depth per cloud) similar to those derived for low- Seyfert 2 galaxies. Following the clumpy dusty torus (Nenkova+08a; Nenkova+08b; Garcia-Bernete+19), the inner radius of the torus () is given as
@@
where is the AGN bolometric luminosity and is the dust sublimation temperature. Assuming the radial extent of the torus () to be about 10 30 times and a dust sublimation temperature of 1500 K, we obtain that the radius of the dusty torus would be in the 3 9 pc range for the derived AGN luminosity of 5.9 1044 erg s-1 (see Table 3). Following the expression by Garcia-Bernete+19 and assuming the typical number of clouds, their opacity, and their radial and angular distribution in Seyfert 2 galaxies (i.e., = 14, =70, = 0 and = 60°, respectively), we derive a torus mass of 0.5 1.5 105 . Thus, the size and mass of the potential torus in GN-z11 would be in the range of those detected in low- Seyfert 2 galaxies with median size and mass of 3.5 3.9 pc and (3.9 5.1) 105 , respectively (Garcia-Bernete+19). Although the SED fitting and colour analysis of the red excess are compatible with the presence of type 2 AGN, the true nature of this potential AGN along with its resemblance to low- counterparts are beyond the scope of this work and needs to await the confirmation of the 10 micron flux and the extension of the photometry into redder wavelengths with additional MIRI imaging.
@@ 4.4.2 Dust emission in compact starbursts
The large star-formation rate and stellar mass along with the compact size of GN-z11 indicate that this galaxy has extreme stellar mass and star formation surface densities. ( = 9.33 102 M⊙ yr-1 kpc-2, Alvarez-marquez+25). In addition, the UV and optical line ratios have shown that this target is a low-metallicity galaxy with low dust extinction (Bunker+23; Alvarez-marquez+25). This is in agreement with scenarios of UV bright galaxies at ¿10 where the narrow line component of the nebular emission is barely extinct as the dust would be evacuated outside these extreme compact starbursts by the outflows (Fiore+23; Ziparo+23; Dekel2023; Li&Dekel+24). In fact, recent studies have identified the existence of dusty outflows in RXCJ2248-ID ( = 6.1; CrespoGomez+25_RXC), and J0217-0208 ( = 6.2; Marques-Chaves+25) as predicted in the Attenuation-Free models (AFM, Ferrara+23; Fiore+23. As GN-z11, these galaxies are also identified as N-emitters (Topping+24; Marques-Chaves+25). These two galaxies show large extinctions in the outflowing ionised gas traced by the broad emission line components identified in their optical spectra. Extinctions of = 1.5 2.5 magnitudes are measured while the blue compact starburst appears almost extinction-free as traced by the narrow Balmer line components and the rest-UV continuum emission. These results show that, despite a low dust-extinction in the UV-bright starburst nucleus, a considerable amount of dust is present in these compact N-emitters. The detection of similar dusty outflows in GN-z11 would require extremely deep spectroscopy with MIRI in order to detect the presence of broad line components in the optical hydrogen and oxygen lines.
@@
Moreover, complex dust and gas structures in the interstellar medium around young compact massive clusters have been detected in low metallicity galaxies. Recent NIRCam and MIRI imaging of the low- galaxy IZw18 (Hirschauer+24) has shown in detail the complexity of the dusty star formation in this extremely metal-poor galaxy (gas-phase oxygen abundance of 3% solar, see Hirschauer+24 and references therein). The presence of massive evolved stars, dust-enshrouded stellar objects as well as shells and filaments of dust/gas due to supernovae has been identified. As shown in the images of IZw18 presented in Hirschauer+24, the dust in and around the stellar clusters could be heated by the strong radiation field, re-radiating in the optical-red and near-infrared wavelength range, and resulting in a flux excess above the stellar+nebular continuum at these wavelengths. Flux excess at wavelengths 0.66 0.86 m as measured in GN-z11 has already been claimed in the low- metal-poor galaxies SBS0335-052E (Reines+08; Adamo+10_SBS) and Haro11 (Adamo+10_Haro11). Notably, Adamo+10_SBS found that many of the young and massive stellar clusters present in SBS0335-052E show flux excesses amounting to 0.2 to 0.5 magnitudes above their predicted stellar continuum at rest-frame 0.8 m. These galaxies have a hard radiation field ( 54.4 eV) as demonstrated by the presence of strong He II 1640 and He II 4686 emission lines (Kehrig+18; Sirressi+22; Wofford+21), and even more recently demonstrated by the detection with MIRI spectroscopy of extended mid-infrared [Ne V] 14.32 m line emission, with an ionization potential of 97 eV, in SBS0335-052E (Mingozzi+25) and IZw18 (Arroyo-Polonio+25).
@@
A photoluminiscence mechanism (Extended Red Emission, or ERE, see Witt+04) has been hypothesised to explain this excess at 0.7 0.9 m in these low-metallicity galaxies (e.g. Reines+08; Adamo+10_SBS). This non-thermal mechanism originates from UV-excited small dust grains or complex molecules in dense circumstellar environments. However, if the excess flux observed in MIRI extends beyond one micron, it would not be consistent with the ERE scenario. If the SED continues to rise, the most likely explanation would be thermal emission of dust close to the sublimation temperature (i.e., 1000 1500 K). Although uncommon, some dense nucleus and young clusters in low metallicity galaxies show near-IR excess compatible with the presence of dust emitting at 600 1100 K (Mrk996, Thuan+08; NGC5253, Vanzi+04; Calzetti+15). The presence of 1000 K dust has been also observed around Wolf-Rayet stars (Marchenko+07; Lau+23), whose existence has been proposed as a potential origin of the unusual large N/O ratio in GN-z11 (Cameron+23; Gunawardhana+25).
@@
Recent NIRSpec and MIRI spectroscopy has clearly detected PAH emission, including the strong 3.3m feature, in the dominant super-cluster (SSC-N) of the nearby galaxy IIZw40 (Lai2025). This indicates that dust and complex molecules are produced and survive in low-metallicity (0.1 0.2 , Leitherer2018) extreme radiation environments as those present in SSC-N, with a radiation field () 8.6 104 times higher than the Habing radiation field ( = 1.6 10-3 erg s-1 cm-2, Habing1968). The radiation field in GN-z11, derived from the extreme UV luminosity ( = 1.71029 erg s-1 Hz-1, Bunker+23) and small size of the galaxy ( = 64 pc, Tacchella+23_GNz11), is a factor 5 higher (i.e., = 4.3 105) than that of SSC-N. Therefore, dust and complex molecules could also be present in the low-metallicity (0.17 , Alvarez-marquez+25) interstellar medium of GN-z11 as in IZw40. The presence and emission processes of dust in very extreme radiation environments such as those present in GN-z11 have to be explored further with MIRI imaging at longer wavelengths covering the near-IR rest-frame.

### v2

@@ 4.3 SF + AGN SED modelling
In addition to the young starburst (Bunker et al., 2023; Tacchella et al., 2023a; Álvarez-Márquez et al., 2025), some studies indicates the presence of high critical density lines and broad emission lines in the ultraviolet suggesting that GN-z11 could host a type 1 active black hole (Maiolino et al., 2024). To evaluate the presence of a potential AGN in GN-z11, we also considered the combination of star formation and an AGN (Type 1 or 2). The potential presence of an AGN would naturally produce a red component associated with the hot dust emission from the torus around an accreting black hole. The best-fit SED models for the different SF+AGN scenarios are presented in the top panel of Fig. 5. Specifically, CIGALE results including type 1 and 2 AGNs are represented in blue and green, respectively, while SYNTHESIZER-AGN result is displayed as a red line. Tables 2 and 3 summarise the photometric and physical results of these models.
@@
Adding a type 1 AGN to the mixed stellar population modelled by CIGALE does not significantly improve the SED fitting beyond 0.66 m. The top and middle panels from Fig. 5 show how the type 1 AGN starts to dominate the emission only above 0.9 m and the best-fit modelled F770W and F1000W fluxes are well below the observed values by factors 1.5 and 2, respectively. The type 1 AGN is modelled in SYNTHESIZER-AGN by a QSO composite template (see Sect. 3.3) that includes the contribution from broad emission lines and the dust emission in the optical-red and near-infrared spectral range. The top and bottom panels in Fig. 5 show that the QSO template is almost flat until rest-frame 1 m and it is not able to reproduce the observed fluxes in the MIRI reddest filters. Only when including an ad hoc extra dusty torus (i.e. highly obscured QSO, see Sect. 3.3) are the modelled F770W and F1000W fluxes (257 and 258 nJy, respectively) compatible with the observed ones. However, the addition of this extra hot dust emission to the empirical spectrum of the QSO is not physically justified. As is shown in Figure 3, the 0.7 m emission from the empirical 1 QSO SED agrees with the spectrum of the = 7.1 luminous QSO J1120+0641, for which JWST/MIRI spectroscopy has revealed a dust emission in the dusty torus with a temperature of about 1400 K (Bosman et al., 2024). This agreement in the SED between the low-intermediate- and high- QSOs makes the physical justification for the presence of such an extra dusty torus emission difficult.
@@
We consider next the presence of a type 2 AGN, whose UV and optical emission is expected to be obscured by the torus. According to the CIGALE best-fit model, the steep optical and near-IR continuum from the type 2 AGN emission increases the SED modelled fluxes up to 254 nJy for the F770W filter, being compatible with the MIRI measured fluxes (see Table 2). Despite the F1000W photometry not being taken into account during the SED fitting due to its low significance, the type 2 scenario yields a F1000W flux (285 nJy) compatible within the uncertainties, whereas the SED best-fit lies below the upper limit for F2100W. Although the available AGN models in CIGALE do not include the contribution from the emission lines produced by the AGN, the analysis carried out indicates that assuming a type 2 AGN yields a better agreement with the observed MIRI fluxes (i.e. optical to red continuum SED slope) than using different type 1 scenarios.
@@
Regardless of the AGN type considered, we derived similar ages for the different stellar populations. Specifically, we estimated ages of 4 6 and 20 40 Myr for the young and mature stellar populations, respectively. Ages for the stellar populations were derived independently for the spatially resolved point-source and extended stellar components ( 11 and 35 Myr, respectively) identified in NIRCam images (Tacchella et al., 2023a). Although the resolution of the MIRI images (i.e. FWHM = 0.27 for F770W) does not allow us to spatially resolve these two components (rest-UV = 0.05 for the extended one), our derived ages are compatible with the previous NIRCam photometry.
@@
To ensure the robustness of these results, we replicated the CIGALE SED fits using the SKIRTOR AGN models (Stalevski et al., 2012, 2016), which include a clumpy dust distribution for the torus. These AGN models yield similar results, whereby only a type 2 AGN can explain the F770W and F1000W fluxes. In the next section, we discuss the general physical scenario drawn from this SED best-fit along with other non-AGN possibilities.
@@ 4.4 Origin of the red excess emission
As has already been shown in the previous sections, the flux in the F770W filter, and in the low-significance (3) F1000W value, shows an excess that cannot be fully explained by the combination of young and intermediate-age stellar populations, even when including the contribution by the nebular emission (i.e. the narrow H line detected with the MRS). This excess hints at the presence of a source radiating at (rest-) wavelengths of 0.66 to 0.86 m. The excess at these wavelengths can be interpreted as the emission produced by either the presence of a dusty torus of an AGN or dust associated with the extremely compact starburst identified in GN-z11.
@@ 4.4.1 Dust in AGN: Dusty torus emission
Although GN-z11 has not been detected in X-rays, a 3 upper limit of (210 keV) ¡ 3 1043 erg s-1 has been derived assuming the typical photon index for narrow-line Seyfert 1 galaxies (Maiolino et al., 2024). In addition, a value of 2.9 1043 erg s-1, based on its H luminosity (Ho et al., 2001a; Álvarez-Márquez et al., 2025), could also be consistent with hosting a type 2 AGN. These upper limits agree well with the predicted (210 keV) based on the SED-fitting assuming the presence of type 1 and 2 AGNs. Using the X-ray extension of the CIGALE models (X-CIGALE, Yang et al. 2020), we obtain (210 keV) = (0.9 0.1) and (2.1 0.3) 1043 erg s-1 for the best-fit SED considering the type 1 and 2 AGNs, respectively. Similar X-ray 210 keV luminosities (i.e. 1.8 1043 and 2.6 1043 erg s-1, respectively) were obtained when applying the (210 keV) relation from Marconi et al. (2004) and the bolometric luminosity derived from the SED fitting (see Table 3). In summary, the predicted X-ray luminosities of the best-fit stellar plus AGN (both type 1 and 2) SEDs are compatible with the non-detection of X-ray emission in GN-z11.
@@
However, the scenario of a type 1 AGN is not favoured by the present results. The potential presence of an undetected, weak broad H line of 2 3 10-18 erg s-1 cm-2 tracing a type 1 AGN (Álvarez-Márquez et al. 2025) would contribute to about 18 27 nJy to the F770W flux, which is not sufficient to explain the excess of 91 nJy found during the SED fitting (see Sect. 4.3). In addition, Figure 3 displays how luminous intermediate 1 (Glikman et al., 2006; Selsing et al., 2016) and 7 QSOs (Bosman et al., 2024) have a SED at rest-wavelengths 0.6 0.8 m flatter than otherwise detected at F770W and F1000W in GN-z11. The SEDs of these QSOs are consistent with hot dust from the black hole surrounding torus emitting at a temperature of about 1400 K, contributing at wavelengths of 1 m and above (e.g. Bosman et al. 2024). Moreover, our SYNTHESIZER-AGN best SED fit combining the stellar populations and type 1 AGN requires an additional ad hoc red source (see Sect 4.3) in addition to the hot ( 1300 1400 K) dust component to explain the observed flux excess at 0.66 and 0.86 m. As was already mentioned in the previous section, this extra component is not physically motivated, and therefore the scenario of a standard type 1 AGN is not supported by the present data.
@@
Under the scenario of a Seyfert 2 contribution, the AGN bolometric luminosity derived by CIGALE can be used to estimate the size and mass of the dusty torus assuming that its physical parameters (i.e. the radial extent and inclination of the torus, the number of clouds along the equator of the torus, the angular and radial distribution of clouds, and the optical depth per cloud) are similar to those derived for low- Seyfert 2 galaxies. Following the clumpy dusty torus (Nenkova et al., 2008a, b; García-Bernete et al., 2019), the inner radius of the torus () is given as
@@
where is the AGN bolometric luminosity and is the dust sublimation temperature. Assuming the radial extent of the torus () to be about 10 30 times and a dust sublimation temperature of 1500 K, we obtain that the radius of the dusty torus would be in the 3 9 pc range for the derived AGN luminosity of 5.9 1044 erg s-1 (see Table 3). Following the expression by García-Bernete et al. (2019) and assuming the typical number of clouds, their opacity, and their radial and angular distribution in Seyfert 2 galaxies (i.e. = 14, = 70, = 0 and = 60°, respectively), we derive a torus mass of 0.5 1.5 105 . Thus, the size and mass of the potential torus in GN-z11 would be in the range of those detected in low- Seyfert 2 galaxies with median size and mass of 3.5 3.9 pc and (3.9 5.1) 105 , respectively (García-Bernete et al., 2019). Although the SED fitting and colour analysis of the red excess are compatible with the presence of a type 2 AGN, the true nature of this potential AGN along with its resemblance to low- counterparts are beyond the scope of this work and need to await the confirmation of the 10 micron flux and the extension of the photometry into redder wavelengths with additional MIRI imaging.
@@ 4.4.2 Dust emission in compact starbursts
The high SFR and stellar mass along with the compact size of GN-z11 indicate that this galaxy has extreme stellar mass and star formation surface densities ( = 9.33 102 M⊙ yr-1 kpc-2, Álvarez-Márquez et al. 2025). In addition, the UV and optical line ratios have shown that this target is a low-metallicity galaxy with low dust extinction (Bunker et al., 2023; Álvarez-Márquez et al., 2025). This is in agreement with scenarios of UV bright galaxies at ¿10 in which the narrow-line component of the nebular emission is barely extinct as the dust would be evacuated outside these extreme compact starbursts by the outflows (Fiore et al., 2023; Ziparo et al., 2023; Dekel et al., 2023; Li et al., 2024). In fact, recent studies have identified the existence of dusty outflows in RXCJ2248-ID ( = 6.1; Crespo Gómez et al. 2025) and J0217-0208 ( = 6.2; Marques-Chaves et al. 2025) as was predicted in the Attenuation-Free model (AFM, Ferrara et al. 2023; Fiore et al. 2023. Similarly to GN-z11, these galaxies are also identified as N-emitters (Topping et al., 2024; Marques-Chaves et al., 2025). These two galaxies show large extinctions in the outflowing ionised gas traced by the broad emission line components identified in their optical spectra. Extinctions of = 1.5 2.5 magnitudes are measured, while the blue compact starburst appears almost extinction-free as traced by the narrow Balmer line components and the rest-UV continuum emission. These results show that, despite a low dust-extinction in the UV-bright starburst nucleus, a considerable amount of dust is present in these compact N-emitters. The detection of similar dusty outflows in GN-z11 would require extremely deep spectroscopy with MIRI in order to detect the presence of broad line components in the optical hydrogen and oxygen lines.
@@
Moreover, complex dust and gas structures in the interstellar medium around young compact massive clusters have been detected in low-metallicity galaxies. Recent NIRCam and MIRI imaging of the low- galaxy IZw18 (Hirschauer et al., 2024) has shown in detail the complexity of the dusty star formation in this extremely metal-poor galaxy (which has a gas-phase oxygen abundance of 3% solar, see Hirschauer et al. 2024 and references therein). The presence of massive evolved stars, dust-enshrouded stellar objects, and shells and filaments of dust and/or gas due to supernovae has been identified. As is shown in the images of IZw18 presented in Hirschauer et al. (2024), the dust in and around the stellar clusters could be heated by the strong radiation field, re-radiating in the optical-red and near-infrared wavelength range, and resulting in a flux excess above the stellar+nebular continuum at these wavelengths. Flux excess at wavelengths of 0.66 0.86 m as measured in GN-z11 has already been claimed in the low- metal-poor galaxies SBS0335-052E (Reines et al., 2008; Adamo et al., 2010b) and Haro11 (Adamo et al., 2010a). Notably, Adamo et al. (2010b) found that many of the young and massive stellar clusters present in SBS0335-052E show flux excesses amounting to 0.2 to 0.5 magnitudes above their predicted stellar continuum at rest-frame 0.8 m. These galaxies have a hard radiation field ( 54.4 eV), as is demonstrated by the presence of strong He II 1640 and He II 4686 emission lines (Kehrig et al., 2018; Sirressi et al., 2022; Wofford et al., 2021), and even more recently demonstrated by the detection with MIRI spectroscopy of extended mid-infrared [Ne V] 14.32 m line emission, with an ionisation potential of 97 eV, in SBS0335-052E (Mingozzi et al., 2025) and IZw18 (Arroyo-Polonio et al., 2025).
@@
A photoluminiscence mechanism (Extended Red Emission, or ERE, see Witt and Vijh 2004) has been hypothesised to explain this excess at 0.7 0.9 m in these low-metallicity galaxies (e.g. Reines et al. 2008; Adamo et al. 2010b). This non-thermal mechanism originates from UV-excited small dust grains or complex molecules in dense circumstellar environments. However, if the excess flux observed in MIRI extends beyond one micron, it would not be consistent with the ERE scenario. If the SED continues to rise, the most likely explanation would be thermal emission of dust close to the sublimation temperature (i.e. 1000 1500 K). Although uncommon, some dense nucleus and young clusters in low-metallicity galaxies show near-IR excess compatible with the presence of dust emitting at 600 1100 K (Mrk996, Thuan et al. 2008; NGC5253, Vanzi and Sauvage 2004; Calzetti et al. 2015). The presence of 1000 K dust has also been observed around WR stars (Marchenko and Moffat, 2007; Lau et al., 2023), whose existence has been proposed as a potential origin of the unusual large N/O ratio in GN-z11 (Cameron et al., 2023; Gunawardhana et al., 2025).
@@
Recent NIRSpec and MIRI spectroscopy has clearly detected PAH emission, including the strong 3.3m feature, in the dominant super-cluster (SSC-N) of the nearby galaxy IIZw40 (Lai et al., 2025). This indicates that dust and complex molecules are produced and survive in low-metallicity (0.1 0.2 , Leitherer et al. 2018) extreme radiation environments such as those present in SSC-N, with a radiation field () 8.6 104 times higher than the Habing radiation field ( = 1.6 10-3 erg s-1 cm-2, Habing 1968). The radiation field in GN-z11, derived from the extreme UV luminosity ( = 1.71029 erg s-1 Hz-1, Bunker et al. 2023) and small size of the galaxy ( = 64 pc, Tacchella et al. 2023a), is a factor of 5 higher (i.e. = 4.3 105) than that of SSC-N. Therefore, dust and complex molecules could also be present in the low-metallicity (0.17 , Álvarez-Márquez et al. 2025) interstellar medium of GN-z11 as in IZw40. The presence and emission processes of dust in very extreme radiation environments such as those present in GN-z11 have to be explored further with MIRI imaging at longer wavelengths covering the near-IR rest-frame.